# Distributed Koopman Operator Learning for Perception and Safe Navigation

### v1

@@ Abstract
This paper presents a unified and scalable framework for predictive and safe autonomous navigation in dynamic transportation environments by integrating model predictive control (MPC) with distributed Koopman operator learning. High-dimensional sensory data are employed to model and forecast the motion of surrounding dynamic obstacles. A consensus-based distributed Koopman learning algorithm enables multiple computational agents or sensing units to collaboratively estimate the Koopman operator without centralized data aggregation, thereby supporting large-scale and communication-efficient learning across a networked system. The learned operator predicts future spatial densities of obstacles, which are subsequently represented through Gaussian mixture models. Their confidence ellipses are approximated by convex polytopes and embedded as linear constraints in the MPC formulation to guarantee safe and collision-free navigation. The proposed approach not only ensures obstacle avoidance but also scales efficiently with the number of sensing or computational nodes, aligning with cooperative perception principles in intelligent transportation system (ITS) applications. Theoretical convergence guarantees and predictive constraint formulations are established, and extensive simulations demonstrate reliable, safe, and computationally efficient navigation performance in complex environments.
@@ I Introduction
The advent of intelligent transportation systems (ITS) has driven increasing interest in autonomous agents that can safely navigate in dynamic, uncertain environments such as urban streets, plazas, and shared pedestrian zones [yan2025selective]. In these settings, safety and reliability become paramount [tordesillas2021faster]. Ground robots or autonomous platforms moving through cluttered, time-varying environments must continuously account for surrounding obstacles, adapt to uncertainty, and rapidly incorporate new sensory information [siegwart2011introduction]. Among control methodologies, model predictive control (MPC) stands out for ITS-related applications by optimizing trajectories while enforcing constraints such as collision avoidance and kinematic limits [rawlings2017model, gratzer2024two, zhang2025mpc].
@@
Avoiding collisions with moving targets such as vehicles, pedestrians, or crowds is a central concern in ITS applications. Considerable work has addressed static obstacles (see e.g., [farina2015application, li2023survey]), but dynamic obstacle avoidance remains more challenging due to unpredictability, interaction effects, and environmental uncertainty [li2023moving, olcay2024dynamic, wei2022moving]. In transportation settings, predictions of pedestrian flows [jiang2025fam, mai2025evaluating] or vehicle trajectories [wang2025risk, ye2025tp] over short horizons become critical. Indeed, those forecasts directly influence the control decisions in a receding-horizon framework. As a result, prediction-aware MPC formulations, where predicted obstacle regions are incorporated as time-varying constraints, have grown in interest in robotics and ITS communities [lindqvist2020nonlinear, zhu2023real, olcay2024dynamic].
@@
In real-world ITS deployments, perception is often achieved through high-dimensional sensing modalities such as cameras, radar, or LiDAR units mounted on vehicles, roadside units, or drones [xu2025lv]. Extracting and modeling dynamic behavior of obstacles from such data is nontrivial, especially when explicit motion models are unknown or unreliable. A natural remedy is to treat the environment as a spatiotemporal density field e.g., occupancy, flow, or intensity maps, and to learn predictive models directly from raw sensor sequences. This shift toward data-driven forecasting enables better handling of complex, non-Newtonian motion patterns (e.g. crowds, group flows) as probabilistic density predictions [hafner2019learning], which aligns well with ITS needs such as trajectory prediction, flow forecasting, and shared autonomous navigation.
@@
The Koopman operator framework forms a powerful backbone for such data-driven modeling. By lifting nonlinear dynamics into a function space, system evolution becomes linear in the space of observables [mezic2013analysis]. This enables use of linear-system tools for nonlinear tasks like prediction, estimation, and control. In practice, finite-dimensional approximations of the Koopman operator are built from time-series data and have shown success in system identification, control, and spatiotemporal forecasting [mezic2013analysis, bevanda2021koopman, otto2021koopman, williams2015data, proctor2016dynamic]. In ITS and robotics domains, Koopman-based methods have been applied to modeling [gu2023deep, tian2025physically], planning [gutow2020koopman], object tracking [comas2021self], robust control [zhang2022robust], and vehicular mobility systems [manzoor2023vehicular, zheng2024koopman].
@@
Nonetheless, most Koopman-learning methods remain centralized [williams2015data, korda2018linear, brunton2016discovering]. Centralized strategies require full access to the entire dataset at one location, an assumption that fails under ITS constraints, such as limited computational capabilities, distributed edge sensors, privacy rules, and bandwidth bottlenecks. Distributed Koopman learning addresses these limitations by allowing agents or computational nodes to compute local operator estimates and coordinate to form a shared global Koopman model. In our framework, high-dimensional sensory snapshots (e.g. partitions of camera images) are processed in parallel across multiple computational nodes (virtual agents). Each agent works on its share of the lifted observables and participates in a consensus-based least-squares scheme to estimate the global Koopman operator without exchanging raw data. This mechanism enhances scalability, reduces communication overhead, and ensures adaptability. The resulting operator can then propagate high-dimensional sensory states forward in time to generate spatial density forecasts of dynamic obstacles.
@@
Here we propose a novel distributed Koopman learning algorithm where each node processes its local sensor data and collaborates in a consensus-based optimization over a communication graph to collectively recover the global Koopman operator. This approach preserves convexity and interpretability of EDMD while granting theoretical convergence to the centralized solution under mild connectivity and step-size conditions. By avoiding nonconvexity, raw data exchange, and high communication demands, it scales efficiently to high-dimensional lifted spaces, making it well suited for ITS-scale environments such as imaging, urban crowd flows, and vehicular sensing. It also complements prior temporal splitting schemes like [azarbahram2025distributed], filling in the orthogonal dimension of data decomposition.
In recent literature, the Koopman operator framework has emerged as a promising paradigm for data-driven modeling and prediction of complex dynamical systems in robotic and transportation domains. In [bueno2025koopman], the Koopman operator was first employed to model and forecast the motion of unknown moving obstacles in a single-agent unmanned aerial vehicle (UAV) navigation problem, providing a predictive mechanism for anticipatory path planning. Building on this idea, [azarbahram2025distributedKoopman] extended the Koopman-based methodology to a cooperative multi-agent setting through distributed and switched MPC, enabling decentralized collision-free navigation of multiple UAVs in dynamic and uncertain environments.
In contrast to these prior works, the present study introduces, for the first time, a Koopman learning architecture that learns the dynamic obstacle model itself in a fully distributed manner. The learned distributed Koopman model is then integrated into an MPC framework for safe navigation in dynamic environments. Predicted obstacle densities are represented as Gaussian mixture models, converted into confidence ellipses, and subsequently approximated by convex polytopes. These polytopic regions are incorporated as linear constraints within the MPC problem, ensuring collision avoidance and maintaining safety margins across the prediction horizon. Crucially, this perception-to-control pipeline is fully distributed where no centralized fusion is required, thus making it suitable for deployment in large-scale, dynamic, sensor-rich environments. This enables real-time, scalable deployment in large-scale, sensor-rich ITS applications, where distributed sensing and computation are essential for robust and safe autonomous operation.
The remainder of the paper is organized as follows. Section II presents the problem formulation and necessary preliminaries; Section III describes the distributed Koopman algorithm and its integration with MPC; Section IV presents simulation results validating the method; and Section V concludes the paper.
@@
Let $\mathcal{G}=(\mathcal{P},\mathcal{E})$ be an undirected graph with node set $\mathcal{P}:=\{1,\dots,p\}$, $p\in\mathbb{N}$, and edge set $\mathcal{E}\subseteq\mathcal{P}\times\mathcal{P}$. An edge $(i,j)\in\mathcal{E}$ implies the reciprocal edge $(j,i)\in\mathcal{E}$. A path is a sequence of nodes linked by edges; the graph is connected if every pair of nodes is joined by some path. For each node $i\in\mathcal{P}$, its neighborhood is $N(i):=\{j\in\mathcal{P}:(i,j)\in\mathcal{E}\}$. The Laplacian matrix associated with graph $G$ is ${L}=[{L}_{ij}]\in\mathbb{R}^{p\times p}$, defined elementwise by
@@ II Preliminaries and Problem Statement
In this section, first the considered robot model is introduced and then the safe navigation control problem is formulated.
@@ II-A Unicycle robot model and feedback linearization
The unicycle robot model is adopted as a representative abstraction of wheeled autonomous ground vehicles widely studied in ITS communities. Its dynamics capture the essential nonholonomic motion constraints of cars, delivery robots, and small-scale urban service platforms, while remaining analytically tractable for control design and verification. This simplified model provides a canonical testbed for developing and validating predictive control strategies in dynamic traffic-like settings, where motion feasibility, safety margins, and real-time trajectory planning are critical. Beyond ground vehicles, the same modeling principles extend naturally to other nonholonomic platforms such as mobile robots and automated guided vehicles used in cooperative transportation and urban mobility networks. The considered kinematics model of the unicycle robot is captured by the nonlinear system
@@
II-B Problem statement
We address the challenge of safe and predictive navigation in dynamic environments through the integration of distributed perception and MPC. Aerial images captured by a downward-facing camera provide high-dimensional spatiotemporal observations of surrounding moving obstacles, such as pedestrians and vehicles. The motion of these obstacles is modeled using a Koopman operator framework, which lifts the underlying nonlinear dynamics into a higher-dimensional linear space, enabling tractable prediction and analysis. To efficiently process the high-dimensional sensory data, a distributed Koopman learning scheme is employed in which each image snapshot is treated through multiple computational nodes. These local computation nodes collaboratively estimate the Koopman operator via a consensus-based optimization procedure, eliminating the need for centralized computation while preserving global model consistency. The learned operator is then used to forecast future obstacle density maps, which are represented as Gaussian mixture models and converted into convex polytopic regions through confidence ellipses. These predicted regions are incorporated into an MPC formulation governing the motion of a unicycle robot, where they act as time-varying linear constraints to ensure collision-free and anticipatory motion planning. The proposed framework thus enables scalable, cooperative, and perception-driven decision-making suitable for real-time navigation in complex, dynamic, and partially observed transportation environments.
@@ III-A Distributed Koopman modeling of dynamic obstacles
The Koopman operator framework provides a linear representation of nonlinear dynamical systems by acting on functions of the state rather than the state itself. Consider the discrete-time nonlinear system
$x_{k+1}$ $=f(x_{k}),\quad x_{k}\in\mathcal{M}\subseteq\mathbb{R}^{q},$ (7)
where the mapping $f:\mathcal{M}\to\mathcal{M}$ is nonlinear and $q$ denotes the state dimension. Instead of analyzing the state trajectory $\{x_{k}\}$, one studies the evolution of observables $\phi:\mathcal{M}\to\mathbb{R}$ belonging to a Hilbert space $\mathcal{H}$ of scalar-valued functions. The infinite-dimensional Koopman operator $\mathcal{K}:\mathcal{H}\to\mathcal{H}$ is introduced as
$(\mathcal{K}\phi)(x)=\phi(f(x)),\quad\forall\phi\in\mathcal{H},\;x\in\mathcal{M}.$ (8)
Although $f$ is nonlinear, the operator $\mathcal{K}$ acts linearly on the space of observables. To approximate its action in finite dimensions, we select $n$ observables and group them into the vector-valued mapping
$\Phi(x)$ $=[\phi_{1}(x),\phi_{2}(x),\dots,\phi_{n}(x)]^{\top}\in\mathbb{R}^{n},$ (9)
which lifts each state $x\in\mathcal{M}$ into an $n$-dimensional feature space [williams2015data]. With a dataset of $N$ samples $\{x_{k},f(x_{k})\}_{k=1}^{N}$, the lifted snapshots are organized as
$X$ $=[\phi(x_{1}),\phi(x_{2}),\dots,\phi(x_{N})]\in\mathbb{R}^{n\times N},$ (10)
$Y$ $=[\phi(f(x_{1})),\phi(f(x_{2})),\dots,\phi(f(x_{N}))]\in\mathbb{R}^{n\times N}.$ (11)
$Y$ $\approx KX.$ (12)
To enable predictive and safe robot navigation, it is essential to model the motion of unknown moving obstacles without explicit dynamics. We adopt a Koopman operator framework that lifts nonlinear dynamics into a linear space using high-dimensional perceptual data (e.g., images or video snapshots) collected over time. These snapshots capture the spatiotemporal distribution of obstacles and serve as the training data for learning the Koopman operator. Due to the high dimensionality of the data, centralized learning becomes computationally infeasible for real-time updates. To address this, we propose a distributed Koopman learning approach, where snapshot data is investigated using multiple computational nodes (virtual agents). Each node processes local lifted observables and collaboratively contributes to a globally consistent Koopman model via consensus updates. This parallel architecture enables scalable, timely learning suitable for integration into real-time obstacle-aware navigation systems.
In the considered setting, each node $i\in\{1,\dots,p\}$ has access only to a local segment of the lifted observables and their forward-time transitions, i.e., $X_{i}\in\mathbb{R}^{n_{i}\times N}$ and $Y_{i}\in\mathbb{R}^{n_{i}\times N}$, where $\sum_{i=1}^{p}n_{i}=n$. Formally, this local partitioning can be expressed globally as
$X=\begin{bmatrix}X_{1}\\ \vdots\\ X_{p}\end{bmatrix},\quad Y=\begin{bmatrix}Y_{1}\\ \vdots\\ Y_{p}\end{bmatrix}$.
@@
This structure reflects practical scenarios where agents collect or process data from disjoint spatial regions or sensor domains, leading to decentralized data ownership. Each agent aims to collaboratively solve the least-squares problem (13) while only accessing its local data and communicating with its neighbors over a connected undirected graph $\mathcal{G}$. The design objective is for each agent $i$ to compute its corresponding block-column of the global Koopman operator $K$, denoted as $K_{i}\in\mathbb{R}^{n\times n_{i}}$, by leveraging distributed optimization and consensus over $\mathcal{G}$, thereby achieving global Koopman operator learning without central data aggregation.
The group of nodes (virtual agents) aims to solve the problem (13) by a distributed algorithm as
@@
2:Initialize arbitrary $K_{i}(0)\in\mathbb{R}^{n\times n_{i}}$, $S_{i}(0)\in\mathbb{R}^{n\times N}$ for all $i\in\mathcal{P}$ according to (15).
@@
and $L$ is the Laplacian matrix of $\mathcal{G}$. The distributed algorithm for Koopman operator learning is summarized in Algorithm 1. Before introducing the main result in Theorem 1, the following lemma is instrumental.
@@
Suppose the communication graph $\mathcal{G}$ is undirect and connected. There exists
@@
Meanwhile, Because the graph $\mathcal{G}$ is connected, $1_{p}^{\top}L=0$. Thus (22) implies $(\bm{Y}-K^{*}\bm{X}-W^{*})\mathbf{L}=0$, where recall $\mathbf{L}$ is defined by (19). Hence it follows from (14b) that

### v2

@@ Abstract
This paper presents a unified and scalable framework for predictive and safe autonomous navigation in dynamic transportation environments by integrating model predictive control (MPC) with distributed Koopman operator learning. High-dimensional sensory data are employed to model and forecast the motion of surrounding dynamic obstacles. A consensus-based distributed Koopman learning algorithm enables multiple computational agents or sensing units to collaboratively estimate the Koopman operator without centralized data aggregation, thereby supporting large-scale and communication-efficient learning across a networked system. The learned operator predicts future spatial densities of obstacles, which are subsequently represented through Gaussian mixture models. Their confidence ellipses are approximated by convex polytopes and embedded as linear constraints in the MPC formulation to guarantee safe and collision-free navigation. The proposed approach not only ensures obstacle avoidance but also scales efficiently with the number of sensing or computational nodes, aligning with cooperative perception principles in autonomous navigation applications. Theoretical convergence guarantees and predictive constraint formulations are established, and extensive simulations demonstrate reliable, safe, and computationally efficient navigation performance in complex environments.
@@ I Introduction
The advent of autonomous navigation systems has driven increasing interest in autonomous agents that can safely navigate in dynamic, uncertain environments such as urban streets, plazas, and shared pedestrian zones [farina2015application]. In these settings, safety and reliability become paramount [tordesillas2021faster]. Ground robots or autonomous platforms moving through cluttered, time-varying environments must continuously account for surrounding obstacles, adapt to uncertainty, and rapidly incorporate new sensory information [siegwart2011introduction]. Among control methodologies, model predictive control (MPC) stands out for autonomous navigation-related applications by optimizing trajectories while enforcing constraints such as collision avoidance and kinematic limits [rawlings2017model].
@@
Avoiding collisions with moving targets such as vehicles, pedestrians, or crowds is a central concern in autonomous navigation applications. Considerable work has addressed static obstacles (see e.g., [farina2015application, li2023survey]), but dynamic obstacle avoidance remains more challenging due to unpredictability, interaction effects, and environmental uncertainty [li2023moving, olcay2024dynamic, wei2022moving]. Indeed, those forecasts directly influence the control decisions in a receding-horizon framework. As a result, prediction-aware MPC formulations, where predicted obstacle regions are incorporated as time-varying constraints, have grown in interest in robotics and autonomous navigation communities [lindqvist2020nonlinear, zhu2023real, olcay2024dynamic].
@@
In real-world autonomous navigation deployments, perception is often achieved through high-dimensional sensing modalities such as cameras, radar, or LiDAR units mounted on vehicles, roadside units, or drones [xu2025lv]. Extracting and modeling dynamic behavior of obstacles from such data is nontrivial, especially when explicit motion models are unknown or unreliable. A natural remedy is to treat the environment as a spatiotemporal density field e.g., occupancy, flow, or intensity maps, and to learn predictive models directly from raw sensor sequences. This shift toward data-driven forecasting enables better handling of complex, non-Newtonian motion patterns (e.g. crowds, group flows) as probabilistic density predictions [hafner2019learning], which aligns well with autonomous navigation needs such as trajectory prediction, flow forecasting, and shared autonomous navigation.
@@
The Koopman operator theory originates from the seminal work of B.O. Koopman in the early 1930s [koopman1931hamiltonian], where it was introduced as an infinite-dimensional linear operator acting on observables of nonlinear dynamical systems. Although initially developed within the context of Hamiltonian mechanics, its broader implications for nonlinear dynamics were recognized much later. Over the past two decades, the Koopman framework has experienced renewed interest, largely due to its systematic development and popularization, establishing its role as a unifying spectral theory for nonlinear systems and data-driven analysis [mezic2005spectral, mezic2013analysis]. For comprehensive historical background and mathematical foundations, we refer the reader to the expository overview in [mezic2021koopman]. The Koopman operator framework forms a powerful backbone for such data-driven modeling. By lifting nonlinear dynamics into a function space, system evolution becomes linear in the space of observables [mezic2013analysis]. This enables use of linear-system tools for nonlinear tasks like prediction, estimation, and control. In practice, finite-dimensional approximations of the Koopman operator are built from time-series data and have shown success in system identification, control, and spatiotemporal forecasting [mezic2013analysis, bevanda2021koopman, otto2021koopman, williams2015data, proctor2016dynamic]. In autonomous navigation and robotics domains, Koopman-based methods have been applied to modeling [gu2023deep, tian2025physically], planning [gutow2020koopman], object tracking [comas2021self], robust control [zhang2022robust], and vehicular mobility systems [manzoor2023vehicular, zheng2024koopman].
@@
Nonetheless, most Koopman-learning methods remain centralized [williams2015data, korda2018linear, brunton2016discovering]. Centralized strategies require full access to the entire dataset at one location, an assumption that fails under autonomous navigation constraints, such as limited computational capabilities, distributed edge sensors, privacy rules, and bandwidth bottlenecks. Distributed Koopman learning addresses these limitations by allowing agents to compute local operator estimates and coordinate to form a shared global Koopman model. In our framework, high-dimensional sensory snapshots (e.g. partitions of camera images) are processed in parallel across multiple agents. Each agent works on its share of the lifted observables and participates in a consensus-based least-squares scheme to estimate the global Koopman operator without exchanging raw data. This mechanism enhances scalability, reduces communication overhead, and ensures adaptability. The resulting operator can then propagate high-dimensional sensory states forward in time to generate spatial density forecasts of dynamic obstacles.
@@
We propose a distributed Koopman learning algorithm in which agents process local sensor data and collaboratively recover a global Koopman operator via consensus-based optimization over a communication graph. The formulation preserves the convexity and interpretability of EDMD and guarantees convergence to the centralized solution under mild connectivity and stepsize conditions, while avoiding nonconvex optimization, raw data exchange, and excessive communication. This enables scalable learning in high-dimensional lifted spaces, making the approach well suited for autonomous navigation scenarios such as imaging-based perception, urban crowd dynamics, and vehicular sensing. The proposed method complements temporal data-splitting approaches such as [azarbahram2025distributed] by addressing an orthogonal, spatially distributed data regime.
The Koopman operator framework has recently gained traction for data-driven modeling and prediction in robotic and navigation systems. In [bueno2025koopman], Koopman-based models were introduced for forecasting unknown moving obstacles in single unmanned aerial vehicle (UAV) navigation. This idea was extended to cooperative multi-UAV settings in [azarbahram2025distributedKoopman] through distributed and switched MPC for decentralized collision-free navigation. In contrast, the present work introduces a fully distributed Koopman learning architecture that learns the dynamic obstacle model itself without centralized fusion and integrates the learned operator into an MPC framework for safe navigation. Predicted obstacle densities are encoded via Gaussian mixtures and conservative polytopic approximations, which are embedded as linear constraints in MPC, enabling scalable and real-time deployment in large-scale, sensor-rich autonomous systems.
The remainder of the paper is organized as follows. The introduction continues with notations and graph definitions; Section II presents the problem formulation and necessary preliminaries; Section III describes the distributed Koopman algorithm and its integration with MPC; Section IV presents simulation results validating the method; and Section V concludes the paper.
@@
Let ${G}=(\mathcal{P},\mathcal{E})$ be an undirected graph with node set $\mathcal{P}:=\{1,\dots,p\}$, $p\in\mathbb{N}$, and edge set $\mathcal{E}\subseteq\mathcal{P}\times\mathcal{P}$. An edge $(i,j)\in\mathcal{E}$ implies the reciprocal edge $(j,i)\in\mathcal{E}$. A path is a sequence of nodes linked by edges; the graph is connected if every pair of nodes is joined by some path. For each node $i\in\mathcal{P}$, its neighborhood is $N(i):=\{j\in\mathcal{P}:(i,j)\in\mathcal{E}\}$. The Laplacian matrix associated with graph $G$ is ${L}=[{L}_{ij}]\in\mathbb{R}^{p\times p}$, defined elementwise by
@@ II Preliminaries and Problem Statement
In this section, we first give some preliminaries and then the safe navigation control problem is formulated.
@@ II-A Unicycle robot model and feedback linearization
The unicycle robot model is adopted as a representative abstraction of wheeled autonomous ground vehicles widely studied in autonomous navigation communities. Its dynamics capture the essential nonholonomic motion constraints of cars, delivery robots, and small-scale urban service platforms, while remaining analytically tractable for control design and verification. This simplified model provides a canonical testbed for developing and validating predictive control strategies in dynamic traffic-like settings, where motion feasibility, safety margins, and real-time trajectory planning are critical. Beyond ground vehicles, the same modeling principles extend naturally to other nonholonomic platforms such as mobile robots and automated guided vehicles used in cooperative transportation and urban mobility networks. The considered kinematics model of the unicycle robot is captured by the nonlinear system
@@
II-B Koopman Operator Preliminaries
The Koopman operator framework provides a linear representation of nonlinear dynamical systems by acting on functions of the state rather than the state itself. Consider the discrete-time system
$x_{k+1}$ $=f(x_{k}),\quad x_{k}\in\mathbb{R}^{q},$ (6)
where the mapping $f$ is nonlinear and $q$ denotes the state dimension. Instead of analyzing the state trajectory $x_{k}$, one studies the evolution of the observable $\Phi$ belonging to a Hilbert space $\mathcal{H}$ of scalar-valued functions. The infinite-dimensional Koopman operator $\mathcal{K}$ is introduced as
$(\mathcal{K}\Phi)(x)=\Phi(f(x)),\quad\forall\Phi\in\mathcal{H}.$ (7)
Although $f$ is nonlinear, the operator $\mathcal{K}$ acts linearly on the space of observables. To approximate its action in finite dimensions, we select $n$ observables and group them into the vector-valued mapping $\bar{\bm{\Phi}}(x):\mathbb{R}^{q}\to\mathbb{R}^{n}$
$\bar{\bm{\Phi}}(x)$ $=[\Phi_{1}(x),\Phi_{2}(x),\dots,\Phi_{n}(x)]^{\top},$ (8)
which lifts each state $x$ into an $n$-dimensional feature space [williams2015data]. With a dataset of $N$ samples $\{x_{k},f(x_{k})\}_{k=1}^{N}$, the lifted snapshots are organized as
$X$ $=[\bar{\bm{\Phi}}(x_{1}),\bar{\bm{\Phi}}(x_{2}),\dots,\bar{\bm{\Phi}}(x_{N})]\in\mathbb{R}^{n\times N},$ (9)
$Y$ $=[\bar{\bm{\Phi}}(f(x_{1})),\bar{\bm{\Phi}}(f(x_{2})),\dots,\bar{\bm{\Phi}}(f(x_{N}))]\in\mathbb{R}^{n\times N}.$ (10)
$Y$ $\approx KX.$ (11)
II-C Problem statement
We address safe and predictive navigation in dynamic environments by integrating distributed perception with MPC. High-dimensional spatiotemporal observations of moving obstacles, obtained from aerial camera snapshots, are modeled using a Koopman operator that lifts nonlinear obstacle dynamics into a linear predictive representation. To handle the data dimensionality, Koopman learning is performed in a distributed manner across multiple computational nodes via consensus, yielding a globally consistent operator without centralized processing. The resulting predictions are used to forecast obstacle density maps, which are approximated by Gaussian mixtures and converted into convex polytopic regions. These regions enter an MPC formulation for a unicycle robot as time-varying linear constraints, enabling anticipatory and collision-free motion planning in complex dynamic environments.
@@ III-A Distributed Koopman modeling of dynamic obstacles
To enable safe and predictive navigation in dynamic environments, we model the motion of unknown obstacles without assuming explicit dynamics. A Koopman operator framework is adopted to lift nonlinear obstacle motion into a linear representation using high-dimensional perceptual data, such as image or video snapshots, which encode the spatiotemporal evolution of the environment.
This setting naturally motivates a distributed learning architecture in two practical scenarios. In the first, a single UAV observes the entire environment through a high-resolution sensor, but the resulting data dimensionality renders centralized Koopman learning computationally infeasible; the learning task is therefore distributed across onboard computational cores, which act as virtual agents processing subsets of the lifted observables. In the second, a team of UAVs with limited, possibly overlapping fields of view collects partial observations of the environment; here, the agents correspond to physical UAVs that communicate over a network to collaboratively learn a unified Koopman model. In both cases, the common challenge is to learn a global Koopman operator from partitioned, high-dimensional data under communication and computational constraints.
Accordingly, each agent $i\in\{1,\dots,p\}$ has access only to a local segment of the raw one-step data,
$\mathcal{X}_{i}=\begin{bmatrix}\bar{x}_{i,1}&\cdots&\bar{x}_{i,N}\end{bmatrix}\in\mathbb{R}^{q_{i}\times N},$
$\mathcal{Y}_{i}=\begin{bmatrix}\bar{y}_{i,1}&\cdots&\bar{y}_{i,N}\end{bmatrix}\in\mathbb{R}^{q_{i}\times N},$
where $\bar{y}_{i,k}$ denotes the forward transition of $\bar{x}_{i,k}$ and $q_{i}$ is the local data dimension. Local lifting maps $\bar{\bm{\Phi}}_{i}:\mathbb{R}^{q_{i}}\to\mathbb{R}^{n_{i}}$ generate the lifted data
$X_{i}=\bar{\bm{\Phi}}_{i}(\mathcal{X}_{i})\in\mathbb{R}^{n_{i}\times N},\quad Y_{i}=\bar{\bm{\Phi}}_{i}(\mathcal{Y}_{i})\in\mathbb{R}^{n_{i}\times N},$
with $\sum_{i=1}^{p}n_{i}=n$, yielding the global structure
$X=\begin{bmatrix}X_{1}\\ \vdots\\ X_{p}\end{bmatrix},\quad Y=\begin{bmatrix}Y_{1}\\ \vdots\\ Y_{p}\end{bmatrix}$,
@@
consistent with (9)–(10). Each agent aims to compute its block-column $K_{i}\in\mathbb{R}^{n\times n_{i}}$ of the global Koopman operator by collaboratively solving the least-squares problem (12) through distributed optimization and consensus over a connected undirected communication graph G{G}, without central data aggregation.
The group of agents aims to solve the problem (12) by a distributed algorithm as
@@
2:Initialize $K_{i}(0)\in\mathbb{R}^{n\times n_{i}}$, $S_{i}(0)\in\mathbb{R}^{n\times N}$ for all $i\in\mathcal{P}$ according to (15).
@@
and $L$ is the Laplacian matrix of G{G}. The distributed algorithm for Koopman operator learning is summarized in Algorithm 1. Before introducing the main result in Theorem 1, the following lemma is instrumental.
@@
Suppose the communication graph G{G} is undirect and connected. There exists
@@
Meanwhile, Because the graph G{G} is connected, $1_{p}^{\top}L=0$. Thus (22) implies $(\bm{Y}-K^{*}\bm{X}-W^{*})\mathbf{L}=0$, where recall $\mathbf{L}$ is defined by (19). Hence it follows from (14b) that
@@
Remark 1.
The stepsize threshold $\alpha_{\max}$ in Theorem 1 is expressed in terms of the eigenvalues of the matrix $\mathbf{M}$, which depends on the global lifted data through $\bm{X}=\operatorname{diag}(X_{1},\dots,X_{p})$ and on the communication graph. As a result, computing $\alpha_{\max}$ exactly requires centralized spectral information and is not directly feasible in a distributed setting. Importantly, the distributed algorithm does not require the exact value of $\alpha_{\max}$: convergence is guaranteed for any stepsize $\alpha$ satisfying $\alpha<\alpha_{\max}$. In practice, such a stepsize can be selected conservatively using distributedly computable bounds. By combining local evaluations of $\|X_{i}\|_{2}$ (aggregated via max-consensus to obtain $\|\bm{X}\|_{2}=\max_{i}\|X_{i}\|_{2}$) with an upper bound on $\lambda_{\max}(L)$. This yields a common stepsize shared by all agents and ensures convergence of Algorithm 1 without centralized coordination.